The National Association for Women in Science asked recent high school grads if they had taken certain science classes. Of those surveyed, 90 said they had taken physics, 138 said they had taken chemistry and 45 said they had taken both. Thirty-two said they had taken neither. How many recent high school grads were surveyed? Show your work.

Answer by ewatrrr(24785)


Hi
The National Association for Women in Science asked recent high school grads
if they had taken certain science classes. Of those surveyed, 90 said they had
taken physics, 138 said they had taken chemistry and 45 said they had taken
both. Thirty-two said they had taken neither.
Only Physics = 90 - 45 = 45
Only chemistry = 138 - 45 = 93
Neither = 32 | 170 recent high school grads were surveyed